OpenStudy (anonymous):

Find all real values of x such that f(x) = 0. f(x) = x3 − x2 − 4x + 4

OpenStudy (anonymous):

0=x^3-x^2-4x+4 group and remove GCF such that x^2(x-1)-4(x-1)=(x-2)(x+2)(x-1)

OpenStudy (anonymous):

x=2,-2,1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

(x-1)(x^2-4)=0 ==>x=1,x=-2,x=2